A stone is thrown upward with an initial speed of 48 ft/s.


A. what will its maximum height be ?

B. where will it be after 1 s ?


1
Expert's answer
2022-02-15T13:35:07-0500

Given:

"v_0=48\:\rm ft/s"

"g=32.17\:\rm ft/s^2"


A. what will its maximum height be?

"h_{\max}=\frac{v_0^2}{2g}=\frac{(48\:\rm ft/s)^2}{2*32.17\:\rm ft/s^2}=35.8\:\rm ft"

B. where will it be after 1 s?

"y=v_0t-gt^2/2=48*1-32.17*1^2/2=31.9\;\rm ft"